OpenStudy (anonymous):

Find the side length of a square with the given area. 49 ft2 A. 2401 ft B. 343 ft C. 49 ft D. 7 ft @Data_LG2

OpenStudy (anonymous):

The area of a square is: \(A = s^2\) Where 's' is the side. So plug in 49: \(49 = s^2\) Now we find the square root of both sides: \(s = \sqrt{49}\)

OpenStudy (anonymous):

So what's the square root of 49?

OpenStudy (anonymous):

I don't understand :(

OpenStudy (anonymous):

Just find the square root of 49.

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

\(\sqrt{49}\)

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Yes, that's correcct.

OpenStudy (anonymous):

Yes, 7 is your answer.

OpenStudy (anonymous):

This is the formula for the Area of a circle: \(A = s^2\) To find the side length, we plug in the Area and solve for 's': \(49 = s^2\) Now we have to get rid of the exponent of 2, and to do that we find the square root of both sides, giving us: \(s = \sqrt{49}\)

OpenStudy (anonymous):

Because if you multiply 7 to itself, you get 49.

OpenStudy (anonymous):

thx